OpenStudy (anonymous):

need help!! 1.A car has a mass of 1.72x103kg. If the force acting on the car is 6.44x103 N to the east,what is the car’s acceleration? Answer in units of m/s2 2.What is the acceleration of a 56 kg block of cement when pulled sideways with a net force of 526 N? Answer in units of m/s2 6.A 1553.8 kg car is traveling at 26.6 m/s when the driver takes his foot off the gas pedal. It takes 4.9 s for the car to slow down to 20 m/s. How large is the net force slowing the car? Answer in units of N

OpenStudy (caozeyuan):

F=ma

OpenStudy (caozeyuan):

m=1720kg, F=6440N=6440kgms^-2, so what is a?

OpenStudy (caozeyuan):

same thing for the rest

OpenStudy (anonymous):

1. You need to use Newton's second law, F=m*a, for this problem. In order to solve for acceleration, divide force by the mass. a=F/m Now, plug in your variables. You're given Force and mass, so this is just putting numbers into an equation. a=(6.44*10^3)/(1.72*10^3) a=3.744 m/s^2 2. Same equation as the first one: F=m*a In order to solve for acceleration, divide force by the mass. a=F/m You're given Force and mass, so put those numbers into the equation. a=526/56 a=9.393 m/s^2 6. First, you need to solve for acceleration, so use the equation Vf=Vi+at (Vf is final velocity, which would be 20 m/s, and Vi is initial velocity, which would be 26.6 m/s) So to solve for acceleration, subtract Vi and then divide by time. Vf=Vi+at Vf-Vi=at (Vf-Vi)/t=a or a=(Vf-Vi)/t Now you can plug in Vf, Vi, and t to find the acceleration. a=(Vf-Vi)/t a=(20-26.6)/4.9 a=-6.6/4.9 a=-1.347 Acceleration is negative because you're slowing the car down. Now, you can use F=ma to solve. F= 1553.8*-1.347 F=-2092.873 N Hope that helps!

OpenStudy (anonymous):

thank you Poeticalto